OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 8x - y = 31 5x + 3y = 23

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

actually i think the solution is (4,1)

OpenStudy (anonymous):

and the check is easy enough: \[8\times 4-1=32-1=31\] \[5\times 4 + 3\times 1=20+3=23\]

OpenStudy (anonymous):

multiply the first equation by 3 to get \[24x-3y=93\] second one is \[5x+3y=23\] add the equations together to get \[29x=113\] \[x=\frac{113}{29}=4\]

OpenStudy (anonymous):

I misstate 116/29 =4 not 29 8x - y = 31 8x-31=y 5x + 3( 8x-31)= 23 5x+24x - 93 = 23 29x = 23 +93 x = 116/29 x=4

OpenStudy (anonymous):

substitute 4 for x in either equation to find y. for example \[8\times 4-y=31\] \[32-y=31\] \[-y=-1\] \[y=1\]

OpenStudy (anonymous):

5(4) + 3y = 23 20 + 3y =23 3y =23 -20 y= 3/3 y =1 i write equation correct , but I caculator wrong